OpenStudy (anonymous):

The number of plasma TV’s produced varies directly as the number of hours a 10 person assembly line works. If 25 TV’s are produced in 2 ½ hours, what is the constant of variation?

OpenStudy (anonymous):

show work please so i understand it

OpenStudy (anonymous):

y = kx 25 = k(2.5) k = 25/2.5 k = 10 The Constant of Variation is 10.